# Representations of a quantum-deformed Lorentz algebra, Clebsch-Gordan map, and Fenchel-Nielsen representation of quantum complex flat connections at level-$\bm{k}$

### v1

@@ 5.1 Periodic states
When deriving the spectral decomposition for $\bm{L}^{(aux)}_{s},\tilde{\bm{L}}^{(aux)}_{s}$ (or $\bm{L}^{(aux)}_{t},\tilde{\bm{L}}^{(aux)}_{t}$), we should instead require $m_{r}+m_{3}+m_{4}\in 2\mathbb{Z}$ (or $m_{r}+m_{2}+m_{3}\in 2\mathbb{Z}$), then the above computation generalizes straight-forwardly, and the result (5.49) is modified by changing the sum to $\sum_{\frac{m_{r}+m_{3}+m_{4}}{2}\in\mathbb{Z}/N\mathbb{Z}}$ (or $\sum_{\frac{m_{r}+m_{2}+m_{3}}{2}\in\mathbb{Z}/N\mathbb{Z}}$).
5.2 Anti-periodic states
The unitary transformation maps between $\mathcal{H}_{+}$ and $\mathcal{H}_{-}$ is given by $e^{i\pi\bm{m}}$:
$f_{\pm}(\mu,m)\in\mathcal{H}_{\pm}\quad\mapsto\quad e^{i\pi{m}}f_{\pm}(\mu,m)\in\mathcal{H}_{\mp}.$ (5.50)
$e^{i\pi\bm{m}}$ commutes with $\bm{y}^{2},\tilde{\bm{y}}^{2}$ but flips signs of $\bm{x},\tilde{\bm{x}}$: $e^{i\pi\bm{m}}\bm{x}e^{-i\pi\bm{m}}=-\bm{x}$ and $e^{i\pi\bm{m}}\tilde{\bm{x}}e^{-i\pi\bm{m}}=-\tilde{\bm{x}}$.
Secondly, given that $\mathcal{H}_{+}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{N}$, for any $f\in\mathcal{H}_{+}$ and for any $\mu\in\mathbb{R}$, $\{f(\mu,m)\}_{m=1}^{N}$ can be seen as a vector in $\mathbb{C}^{N}$, and it has the discrete Fourier transform
$f\left(\mu,m\right)=\frac{1}{\sqrt{N}}\sum_{n=0}^{N-1}f_{n}(\mu)e^{\frac{2\pi i}{N}mn}.$ (5.51)
We define a unitary map $O:\mathcal{H}_{+}\to\mathcal{H}_{+}$ by
$f_{n}(\mu)\mapsto(-1)^{n}f_{n}(\mu),\qquad\forall\ n=0,\cdots,N-1.$ (5.52)
The unitary map may also be written formally as $\bm{O}f(\mu,m)=f(\mu,m+N/2)$. $O$ is unitary since $\langle Of,Of^{\prime}\rangle=\sum_{n=0}^{N-1}\int\mathrm{d}\mu f_{n}(\mu)^{*}f_{n}^{\prime}(\mu)=\langle f,f^{\prime}\rangle$. $O$ commutes with $\bm{x},\tilde{\bm{x}}$ but flips signs of $\bm{y}^{2},\tilde{\bm{y}}^{2}$: $O\bm{y}^{2}O^{-1}=-\bm{y}^{2}$ and $O\tilde{\bm{y}}^{2}O^{-1}=-\tilde{\bm{y}}^{2}$. $O$ shows that the two representations of Weyl algebra $\bm{x}\bm{y}^{2}=\mathbf{q}^{2}\bm{y}^{2}\bm{x}$ on $L^{2}(\mathbb{R})\otimes\mathbb{C}^{N}$ with flipping $\bm{y}^{2}\to-\bm{y}^{2}$ are unitary equivalent.
By composing these unitary maps, we have $Oe^{i\pi\bm{m}}:\mathcal{H}_{-}\to\mathcal{H}_{+}$, and since $\bm{L},\tilde{\bm{L}}$ leave $\mathcal{H}_{\pm}$ invariant, we have
$Oe^{i\pi\bm{m}}\bm{L}e^{-i\pi\bm{m}}O^{-1}=-\bm{L}\qquad Oe^{i\pi\bm{m}}\tilde{\bm{L}}e^{-i\pi\bm{m}}O^{-1}=-\tilde{\bm{L}}.$ (5.53)
where the right-hand sides acting on $\mathcal{H}_{+}$. This shows that the spectral decomposition of $\bm{L},\tilde{\bm{L}}$ on $\mathcal{H}_{-}$ is unitary equivallent to the spectral decomposition of $-\bm{L},-\tilde{\bm{L}}$ on $\mathcal{H}_{+}$. Therefore, we obtain the same direction integral decomposition as (5.49) for $\mathcal{H}_{-}$
$\mathcal{H}_{-}\simeq\sum_{m_{r}^{\prime}/2\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\rho(\mu_{r},m_{r}^{\prime})^{-1}\,\mathcal{H}_{\mu_{r},m_{r}^{\prime}}.$ (5.54)
But the representation of $\bm{L},\tilde{\bm{L}}$ is different from (5.48) by a minus sign
$\bm{L}f_{-}(\mu_{r},m_{r}^{\prime})=-\ell(r^{\prime})f_{-}(\mu_{r},m_{r}^{\prime}),\qquad\tilde{\bm{L}}f_{-}(\mu_{r},m_{r}^{\prime})=-\ell(r^{\prime})^{*}f_{-}(\mu_{r},m_{r}^{\prime}).$ (5.55)
where $r^{\prime}=\exp[\frac{\pi i}{N}(-ib\mu_{r}-m_{r}^{\prime})]$. In these two formulae, $m_{r}^{\prime}/2\in\mathbb{Z}/N\mathbb{Z}$, i.e. $m_{r}^{\prime}$ is even. We may define the shift $m_{r}=m_{r}^{\prime}+N$, where $m_{r}$ is odd, and $F_{-}(\mu_{r},m_{r})=f_{-}(\mu_{r},m_{r}^{\prime})$, then
$\bm{L}F_{-}(\mu_{r},m_{r})=\ell(r)F_{-}(\mu_{r},m_{r}),\qquad\tilde{\bm{L}}F_{-}(\mu_{r},m_{r})=\ell(r)^{*}F_{-}(\mu_{r},m_{r}).$ (5.56)
where $r=\exp[\frac{\pi i}{N}(-ib\mu_{r}-m_{r})]$. By relabeling $\mathcal{H}_{\mu_{r},m_{r}^{\prime}}\simeq\mathcal{H}_{\mu_{r},m_{r}}$, (5.54) can be written as
$\mathcal{H}_{-}\simeq\sum_{(m_{r}+N)/2\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\rho(\mu_{r},m_{r})^{-1}\,\mathcal{H}_{\mu_{r},m_{r}}.$ (5.57)
From the perspective of eigenstates $\alpha_{r}$, we find for odd $m_{r}$,
$Oe^{i\pi\bm{m}}\alpha_{r}$ $=$ $e^{i\pi(m+N/2)}e^{-\frac{i\pi}{N}m_{r}\frac{N}{2}}e^{-\frac{2\pi i}{k}\left(-\mu_{r}\mu+m_{r}m\right)}\frac{\varphi\left(-{\bf q}^{-1}y^{2}r^{\prime},-\tilde{{\bf q}}^{-1}\tilde{y}^{2}\tilde{r}^{\prime}\right)}{\varphi\left(-{\bf q}y^{-2}r^{\prime},-\tilde{{\bf q}}\tilde{y}^{-2}\tilde{r}^{\prime}\right)}$ (5.58)
$=$ $e^{i\pi\frac{m_{r}^{\prime}}{2}}e^{-\frac{2\pi i}{k}\left(-\mu_{r}\mu+m^{\prime}_{r}m\right)}\frac{\varphi\left(-{\bf q}^{-1}y^{2}r^{\prime},-\tilde{{\bf q}}^{-1}\tilde{y}^{2}\tilde{r}^{\prime}\right)}{\varphi\left(-{\bf q}y^{-2}r^{\prime},-\tilde{{\bf q}}\tilde{y}^{-2}\tilde{r}^{\prime}\right)}=(-1)^{m_{r}^{\prime}/2}\alpha_{r^{\prime}}$
where $m_{r}^{\prime}=m_{r}-N$ is even and $r^{\prime}=\exp[\frac{\pi i}{N}(-ib\mu_{r}-m_{r}^{\prime})]$, $\tilde{r}^{\prime}=\exp[\frac{\pi i}{N}(-ib^{-1}\mu_{r}+m_{r}^{\prime})]$. The unitary transformation maps $\alpha_{r}$ with odd $m_{r}$ to $\alpha_{r^{\prime}}$ (up to a constant phase) with even $m_{r}^{\prime}$. Then all the results in Section 5.1 follows for $\alpha_{r^{\prime}}$.
Combining (5.49) and (5.57), we obtain the spectral decompostion of the full Hilbert space
$\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}\simeq\sum_{m_{r}\in\mathbb{Z}/k\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\rho(\mu_{r},m_{r})^{-1}\,\mathcal{H}_{\mu_{r},m_{r}}.$ (5.59)
containing both even and odd $m_{r}$. This completes the proof of Theorem 5.1.
@@ 6 Changing triangulation
$i_{3}\left(\bm{Z}_{1}\right)$ $=$ $-\bm{X}+i\pi,\qquad i_{3}\left(\bm{Z}_{2}\right)\to-2L_{3}+2\bm{Y}+i\pi,$
$i_{3}\left(\bm{Z}_{3}\right)$ $=$ $-L_{1}+L_{2}-L_{3}-L_{4}-\bm{X}+i\pi,$
$i_{3}\left(\bm{Z}_{4}\right)$ $=$ $2L_{3}+2L_{4}+\bm{X}-2\bm{Y}+i\pi,\qquad i_{3}\left(\bm{Z}_{5}\right)=L_{1}-L_{2}+L_{3}+L_{4}+\bm{X}-2\bm{Y}+i\pi,$
$i_{3}\left(\bm{Z}_{6}\right)$ $=$ $-L_{1}-L_{2}-L_{3}-L_{4}-2\bm{Y}+i\pi$ (6.13)
@@
$i_{3}\left(\bm{z}_{1}\right)$ $=$ $-\bm{x}^{-1},\qquad i_{3}\left(\bm{z}_{2}\right)=-{\lambda_{3}^{-2}}\bm{y}^{2},\qquad i_{3}\left(\bm{z}_{3}\right)=-\lambda_{1}^{-1}\lambda_{2}\lambda_{3}^{-1}\lambda_{4}^{-1}\bm{x}^{-1},$
$i_{3}\left(\bm{z}_{4}\right)$ $=$ $-\mathbf{q}{\lambda_{3}^{2}\lambda_{4}^{2}}\bm{x}\bm{y}^{-2},\qquad i_{3}\left(\bm{z}_{5}\right)=-\mathbf{q}\lambda_{2}^{-1}\lambda_{1}\lambda_{3}\lambda_{4}\bm{x}\bm{y}^{-2},\qquad i_{3}\left(\bm{z}_{6}\right)=-{(\lambda_{1}\lambda_{2}\lambda_{3}\lambda_{4})^{-1}}\bm{y}^{2}.$ (6.16)
@@
If we denote by $V_{s}$ the unitary transformation from $\mathcal{H}$ to the direct integral representation in (5.59) for the S-cycle trace, the composition $V_{s}\circ\phi(-\bm{z}_{3},-\tilde{\bm{z}}_{3})^{-1}$ maps from $\mathcal{H}^{\prime}$ to the direct integral representation.
@@ Acknowledgements
M.H. acknowledges Chen-Hung Hsiao and Qiaoyin Pan for helpful discussions. M.H. receives supports from the National Science Foundation through grant PHY-2207763 and the College of Science Research Fellowship at Florida Atlantic University. M.H. also receives support from the visiting professorship at FAU Erlangen-Nürnberg at the early stage of this work.

### v2

@@ 5 Spectral decomposition
[…]
$\mathrm{d}P_{\lambda,\tilde{\lambda}}=\sum_{m_{r}^{\prime}\in\mathbb{Z}/N\mathbb{Z}}\mathrm{d}\mu_{r}\mathrm{d}m_{r}\,\Theta(\mu_{r})\delta(m_{r}-m^{\prime}_{r})\varrho(\mu_{r},m_{r})^{-1}\mathcal{S}_{\lambda_{1}}^{-1}\mathcal{D}_{\lambda_{2}}{\cal F}^{-1}|\psi_{r}\rangle\langle\psi_{r}|{\cal F}\mathcal{D}_{\lambda_{2}}^{-1}\mathcal{S}_{\lambda_{1}}.$ (5.33)
Given that $|\psi_{r}\rangle$ diagonalizing $(\bm{L},\tilde{\bm{L}})$, $\mathfrak{U}_{s}^{-1}|\psi_{r}\rangle$ and $\mathfrak{U}_{t}^{-1}|\psi_{r}\rangle$ diagonalize $(\bm{L}_{s},\tilde{\bm{L}}_{s})$ and $(\bm{L}_{t},\tilde{\bm{L}}_{t})$ respectively. Therefore, for any $f\in\mathcal{H}$,
$f_{s}(\mu_{r},m_{r})=\langle\psi_{r}|\mathfrak{U}_{s}|f\rangle,\qquad f_{t}(\mu_{r},m_{r})=\langle\psi_{r}|\mathfrak{U}_{t}|f\rangle,$ (5.34)
satisfy
$\langle f\mid f^{\prime}\rangle$ $=$ $\sum_{m_{r}\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\varrho(\mu_{r},m_{r})^{-1}f_{s}(\mu_{r},m_{r})^{*}f^{\prime}_{s}(\mu_{r},m_{r})$ (5.35)
$=$ $\sum_{m_{r}\in\mathbb{Z}/N\mathbb{Z}}\int_{0}^{\infty}\mathrm{d}\mu_{r}\,\varrho(\mu_{r},m_{r})^{-1}f_{t}(\mu_{r},m_{r})^{*}f^{\prime}_{t}(\mu_{r},m_{r}),$ (5.36)
and
$\bm{L}_{s}f_{s}(\mu_{r},m_{r})$ $=$ $\ell(r)f_{s}(\mu_{r},m_{r}),\qquad\tilde{\bm{L}}_{s}f_{s}(\mu_{r},m_{r})=\ell(r)^{*}f_{s}(\mu_{r},m_{r}),$ (5.37)
$\bm{L}_{t}f_{t}(\mu_{r},m_{r})$ $=$ $\ell(r)f_{t}(\mu_{r},m_{r}),\qquad\tilde{\bm{L}}_{t}f_{t}(\mu_{r},m_{r})=\ell(r)^{*}f_{t}(\mu_{r},m_{r}).$ (5.38)
Representing states in $\mathcal{H}$ by functions $f_{s}(\mu_{r},m_{r})$ or $f_{t}(\mu_{r},m_{r})$ define the FN representation with respect to the S-cycle or T-cycle respectively.
@@ 6 Changing triangulation
$i_{3}\left(\bm{Z}_{1}\right)$ $=$ $-\bm{U},\qquad i_{3}\left(\bm{Z}_{2}\right)\to-2L_{3}+\bm{Y}+i\pi,$
$i_{3}\left(\bm{Z}_{3}\right)$ $=$ $-L_{1}+L_{2}-L_{3}-L_{4}-\bm{U},$
$i_{3}\left(\bm{Z}_{4}\right)$ $=$ $2L_{3}+2L_{4}+\bm{U}-\bm{Y}+2i\pi,\qquad i_{3}\left(\bm{Z}_{5}\right)=L_{1}-L_{2}+L_{3}+L_{4}+\bm{U}-\bm{Y}+2i\pi,$
$i_{3}\left(\bm{Z}_{6}\right)$ $=$ $-L_{1}-L_{2}-L_{3}-L_{4}+\bm{Y}+i\pi$ (6.13)
@@
$i_{3}\left(\bm{z}_{1}\right)$ $=$ $\bm{u}^{-1},\qquad i_{3}\left(\bm{z}_{2}\right)=-{\lambda_{3}^{-2}}\bm{y},\qquad i_{3}\left(\bm{z}_{3}\right)=\lambda_{1}^{-1}\lambda_{2}\lambda_{3}^{-1}\lambda_{4}^{-1}\bm{u}^{-1},$
$i_{3}\left(\bm{z}_{4}\right)$ $=$ $\mathbf{q}{\lambda_{3}^{2}\lambda_{4}^{2}}\bm{u}\bm{y}^{-1},\qquad i_{3}\left(\bm{z}_{5}\right)=\mathbf{q}\lambda_{2}^{-1}\lambda_{1}\lambda_{3}\lambda_{4}\bm{u}\bm{y}^{-1},\qquad i_{3}\left(\bm{z}_{6}\right)=-{(\lambda_{1}\lambda_{2}\lambda_{3}\lambda_{4})^{-1}}\bm{y}.$ (6.16)
@@
If we denote by $V_{s}$ the unitary transformation from $\mathcal{H}$ to the direct integral representation for the S-cycle trace, the composition $V_{s}\circ\phi(-\bm{z}_{3},-\tilde{\bm{z}}_{3})^{-1}$ maps from $\mathcal{H}^{\prime}$ to the direct integral representation.
@@ Acknowledgements
M.H. acknowledges Chen-Hung Hsiao and Qiaoyin Pan for helpful discussions. M.H. receives supports from the National Science Foundation through grant PHY-2207763, the College of Science Research Fellowship at Florida Atlantic University, and the Blaumann Foundation. M.H. also receives support from the visiting professorship at FAU Erlangen-Nürnberg at the early stage of this work.
@@ Appendix A Quantum dilogarithm
For $N=1$, $\gamma(x,n)=\gamma(x,0)\equiv\gamma(x)$ is Faddeev’s quantum dilogarithm in e.g. [7, 27]:
$\gamma(x)=\exp\left[\frac{1}{4}\int_{\mathbb{R}+i0^{+}}\frac{\mathrm{d}w}{w}\frac{e^{2iwx}}{\sinh\left(b^{-1}{w}\right)\sinh(bw)}\right]$ (A.10)
@@ Appendix B Fock-Goncharov coordinate and holonomies
Let us consider the ideal triangulation in FIG.1. $\gamma_{i}$ is the loop travels around the $i$-th hole counter-clockwisely. All $\gamma_{i}$ share the same base point represented by a snake pointing from the 1st hole to the 2nd hole along the edge 6 with the fin inside the triangle with vertices 1,2,4. We obtain
$H(\gamma_{1})$ $=$ $E\left(Z_{6}\right).\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right)E\left(Z_{4}\right).\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right),$ (B.70)
$H(\gamma_{2})$ $=$ $\left(\begin{array}[]{cc}0&1\\ -1&0\\ \end{array}\right)^{-1}\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right)E\left(Z_{3}\right)\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right)E\left(Z_{5}\right)\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right)E\left(Z_{1}\right)\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right)E\left(Z_{6}\right)\left(\begin{array}[]{cc}0&1\\ -1&0\\ \end{array}\right),$ (B.83)
$H(\gamma_{3})$ $=$ $\mathbf{M}_{3}^{-1}E\left(Z_{5}\right)\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right)E\left(Z_{2}\right)\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right)\mathbf{M}_{3},$ (B.88)
$H(\gamma_{4})$ $=$ $\mathbf{M}_{4}^{-1}E\left(Z_{2}\right)\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right)E\left(Z_{3}\right)\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right)E\left(Z_{4}\right)\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right)E\left(Z_{1}\right)\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right)\mathbf{M}_{4},$ (B.97)
where
$\mathbf{M}_{3}$ $=$ $\left(\begin{array}[]{cc}0&1\\ -1&0\\ \end{array}\right)^{-1}\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right)E\left(Z_{1}\right)\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right)E\left(Z_{6}\right)\left(\begin{array}[]{cc}0&1\\ -1&0\\ \end{array}\right),$ (B.106)
$\mathbf{M}_{4}$ $=$ $\left(\begin{array}[]{cc}0&1\\ -1&0\\ \end{array}\right)^{-1}\left(\begin{array}[]{cc}1&0\\ 1&1\\ \end{array}\right)\mathbf{M}_{3}.$ (B.111)
The set of $H(\gamma_{i})$ satisfies
$H(\gamma_{1})H(\gamma_{2})H(\gamma_{3})H(\gamma_{4})=1.$ (B.112)